OpenStudy (anonymous):

Please draw a graph of the circle with the equation (x+1)^2+(y+2)^2=17

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

The equation for a circle is of the form: (x-h)^2 + (y-k)^2= r^2 where r is the radius.

OpenStudy (anonymous):

Can you give four or five points on this circle. i see 0,2 and 0,-6

OpenStudy (anonymous):

(h,k) are center points. x+1=0 x=-1 y+2=0 y=-2 So, (-1,-2) are the center points. \[\sqrt{17}=r\] will give you the radius. or the distance from the center to the circumference.

OpenStudy (anonymous):

3,-3? -5,-3?

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

yes! That is correct!! :D

OpenStudy (anonymous):

Thank you soooo much !!!